RC meeting and Run Plan, Tuesday 2nd April
RC Meeting Information
RC Daily Meetings are at 3:30 PM, 2nd floor Counting House Meeting Room
- Bluejeans remote connection information:
To join the Meeting: https://bluejeans.com/861439920 To join via Room System: Video Conferencing System: bjn.vc -or-188.8.131.52 Meeting ID : 861439920 To join via phone : 1) Dial: +1.408.740.7256 (US (San Jose)) +1.888.240.2560 (US Toll Free)) +1.408.317.9253 (US (Primary, San Jose)) (see all numbers - http://bluejeans.com/numbers) 2) Enter Conference ID : 861439920
- KaonLT-SIDIS Group 2018-2019 Shift Schedule
- Hall-C Controls
- Hall-C Logbook
- MCC White Board
- KaonLT Status Page
- Currently running the Q2 = 3.0, W = 3.14, x = 0.25, centre (SHMS = 6.91 degrees) setting
- Statistics coming in nicely, rate dependency test (running at 40uA for a little) completed
- Plan is to just stick at 60uA now
- Rotation issue over the weekend, resolved, see the logentry from yesterday - https://logbooks.jlab.org/entry/3674507 and the new notice below!
- Remote rotation is now allowed again but PLEASE pay attention and follow the procedure when rotating
- Short downtime today whilst they try to save some RF cavities from themselves
Current Run Plan
- We stay at this setting until Thursday day shift when we switch to the left setting (SHMS = 9.91 degrees, same momenta)
- We will run the aluminium dummy target for 3 hours overnight tonight and then switch back to LH2 for the remaining time
- Barring any major incidents, on track to hit the stats goal for this setting
Notices in Effect for Shift Crew
- Please take a screenshot of the spectrometer angles when they are moved to a new position and make a log entry. If working on a kinematic that lasts several days take an image daily.
- Pay attention to the notices posted in the counting house regarding HMS Q1/Q2.
- If HMS interlocks when ramping, check the live camera image of the HMS power supplies for smoke/water before clearing the interlock.
- Hit enter once you type in the angle when rotating(!)